# LVLM-Count: Enhancing the Counting Ability of Large Vision-Language Models

### v3

@@ Abstract
Counting is a fundamental operation for various real-world visual tasks, requiring both object recognition and robust counting capabilities. Despite their advanced visual perception, large vision-language models (LVLMs) are known to struggle with counting tasks. In this work, we evaluate the performance of several recent LVLMs on visual counting tasks across multiple counting and vision datasets. We observe that while their performance may be less prone to error for small numbers of objects, they exhibit significant weaknesses as the number of objects increases. To alleviate this issue, we propose a simple yet effective baseline method that enhances LVLMs’ counting ability for large numbers of objects using a divide-and-conquer approach. Our method decomposes counting problems into sub-tasks. Moreover, it incorporates a mechanism to prevent objects from being split during division, which could otherwise lead to repetitive counting—a common issue in a naive divide-and-conquer implementation. We demonstrate the effectiveness of this approach across various datasets and benchmarks, establishing it as a valuable reference for evaluating future solutions.
@@ 1 Introduction
Recent advancements in the field of large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), coupled with the massive scale of web-scraped data used to train them, have resulted in unprecedented zero-shot capabilities across diverse tasks. However, evaluations of LVLMs’ capabilities have revealed weaknesses in numerical reasoning Yin et al. (2023); Xu et al. (2024); Yang et al. (2023).
In this work, we focus specifically on the visual counting task, one of the most fundamental aspects of numerical reasoning. We observe that although LVLMs perform well in counting small numbers of objects—typically fewer than 20—their accuracy deteriorates with larger quantities. Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al. (2017) we propose a simple yet effective baseline method to alleviate this issue. Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into sub-images, and prompts the LVLM to count the objects of interest in each sub-image. The counts from the sub-images are then aggregated to make the final prediction. A key feature of our pipeline is a mechanism that prevents objects of interest from being split by the dividing lines, which could otherwise lead to double-counting. To have this mechanism two existing pre-trained detection and segmentation models are employed in the pipeline. The workflow is illustrated in Figure 1.
@@
    We propose a solution for object-aware division. Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure 2(a)). The proposed solution divides images without cutting through objects of interest specified by an arbitrary prompt.
@@
Figure 2: Comparison of the naive and the object-aware division. The objects of interest are the circles. In Figure 2(a), we illustrate a naive division of the input image, which is divided into equally sized sub-images with straight lines. In Figure 2(b), we illustrate the object-aware division, which avoids cutting through circles. In Figure 2(c), we illustrate the counting error of GPT-4o for images with randomly positioned circles. The absolute counting error is the absolute difference between the ground truth and the number predicted by GPT-4o. The results are averaged over three trials.
@@ 2 Related Work
Text-based counting models. With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world models have been proposed that are trained specifically for counting. Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting models by Amini-Naieni et al. (2023); Dai et al. (2024); Kang et al. (2024); Amini-Naieni et al. (2024) have started to demonstrate comparable or superior accuracy. In addition, Shi et al. (2024) introduce TFOC, a counting model that does not require any counting-specific training. Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation masks that determine the output number.
@@ 3.1 Area Detection
In this part of the pipeline, we assume that we are given a counting question $Q$ along with an image. The question $Q$ contains an expression $E$ that specifies a set of objects of interest. The expression $E$ distinguishes these objects from objects of other categories or the same category but with different attributes present in the image. By employing an LLM, the expression $E$ is extracted from $Q$. For example, let $Q$ be “How many brown eggs are in the image?”. $Q$ is given to an LLM, which is prompted to return the expression $E$, “brown eggs”, referring to the objects we want to count. After $E$ is extracted, it is provided as input to GroundingDINO along with the image. The output of GroundingDINO may be a single bounding box or a set of bounding boxes that have relevance to $E$ beyond a certain threshold. These bounding boxes often overlap and typically contain repeated objects. Thus, all the overlapping output bounding boxes are merged. After merging, a set of non-overlapping areas of interest may remain. We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage. This process is illustrated in Figure 3.
@@ 3.2 Target Segmentation
The cropped images from the first stage contain objects of interest, and the ultimate goal is to divide them without cutting through those objects. However, a prerequisite for implementing such a mechanism is to first detect and localize the objects of interest. Each cropped image is fed into an open-world detection model along with $E$. The output of the open-world detection model produces a bounding box for each object of interest. The bounding boxes are then given as input to a segmentation model, which returns segmentation masks for the objects within each bounding box. We illustrate an example of this process in Figure 4.
@@
Figure 4: Illustration of the target segmentation step of LVLM-Count. The goal is to produce all the instance masks for $E$ set to “brown egg”. The cropped detected area from Figure 3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure 4(a). Figure 4(a) is then given as input to SAM, which produces the output shown in Figure 4(b).
@@
How to determine the segmentation masks. We employ a pre-trained segmentation model, specifically SAM (Kirillov et al., 2023), for the segmentation task. This model accepts bounding boxes as prompts and generates masks covering the most prominent objects within these boxes. However, we do not use SAM’s output masks directly, as in crowded scenes or cases with multiple occlusions, the masks often overlap, making it difficult to identify reliable division paths. To address this issue, we apply several post-processing steps. First, we perform non-maximum suppression on SAM-generated masks to eliminate those with significant overlaps corresponding to less certain bounding boxes. Additionally, we apply an erosion function to the segmentations, ensuring adjacent masks maintain a minimum separation of two pixels. For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix C. The final processed masks for each cropped area are then passed to the next stage of our pipeline.
Robustness to the Accuracy of Area Detection and Target Segmentation Stages.
Although we employ GroundingDINO—a state-of-the-art model for grounding and detection—it lacks the flexibility and robustness of LVLMs when generalizing to unseen data and complex concepts. Consequently, we minimizes dependence on the accuracy of the detection model used for area detection and target segmentation. Rather than prioritizing accuracy, we emphasize: i) not missing any region containing objects of interest during area detection, and ii) not missing segmentation masks for any objects of interest during target segmentation. These objectives are easily achieved by setting GroundingDINO’s detection threshold to a very low value in both stages.
@@
We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section 4.2. This dataset is particularly difficult due to frequent occlusion, camouflage, and complex backgrounds Arteta et al. (2016). Furthermore, in our ablation studies, we test a variant of our pipeline that removes the detection model entirely. Instead, SAM is configured to generate segmentation masks for all objects in the scene. As shown in Table 5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
@@ 3.3 Object-aware Division
How to determine the starting and ending points of the paths. We utilize an unsupervised and non-parametric approach, to obtain the start and end points of the paths. A few pixels are sampled from each of the masks. To determine the location of the division paths on the $x$-axis, the samples taken from the masks are projected onto the $x$-axis. The projected points are automatically clustered using a non-parametric mean-shift algorithm. Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a division path. In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known. In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively. Note that using this technique, we obtain the appropriate coordinates for the division paths, as well as the number of paths. For example, if there is only one cluster along the $x$-axis, no division is required, and if there are two clusters, one vertical path will divide the image into two parts. We illustrate this approach in Figure 5.
@@
How to draw the paths. Previous step obtains the endpoints of the division paths. Assume $P_{s}=(x,0)$ and $P_{e}=(x,h)$ are the start and end points of a vertical division path, respectively. In an ideal case where there are no masks in the path of a straight line connecting the two points, this path will be drawn by connecting all the pixels on the straight line. However, there are potential masks that can be considered obstacles blocking the path. In other words, beginning from $P_{s}$, the line needs to go around these obstacles to reach $P_{e}$. Consequently, we treat this as a 2-dimensional path-finding problem. To solve the problem, we build a 2D binary map, $I_{B}$, where the pixels covered by the masks are turned into black, indicating them as obstacles, and all the other pixels are turned into white, showing they are open for passage. This binary image $I_{B}$ is mapped into a graph $G$, where each white pixel is a node, and it is connected to all of its white neighboring pixels. We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance. The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure 6. The path-finding algorithm is run for all division coordinates. Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image.
@@ 4 Experiments
In this section, we present the performance results of different LVLMs and their enhanced performance results using our method on a counting-specific dataset, a counting benchmark taken from a popular vision datasets, and a challenging counting benchmark that we propose using emoji icons. Additionally, we show the success of our design in enhancing counting performance on one of the most challenging counting datasets that features heavy occlusions and complex backgrounds. The code to reproduce the experiments can be found here: https://github.com/mrghofrani/lvlm-count.
@@ 4.1 Datasets and Benchmarks
Emoji-Count. To our knowledge, no counting benchmark exists with large number of objects in a scene involving complex reasoning. To this end, we propose a challenging counting benchmark using emoji icons. From the $1816$ standard emoji icons, we remove those that directly overlap with concepts demonstrated by other icons. We then group the remaining $1197$ icons into $82$ classes. In each class, there are icons from the same or similar object categories, but with subtle differences that require complex reasoning to distinguish. For each of the $82$ classes, an empty $1024\times 1024$ image is first created. This image is filled with up to six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image. For each image, we create questions that ask the number of instances of the available categories in the image. This results in 415 image-question pairs. We illustrate two examples of this dataset in Figure 7.
@@
Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure 7(a), the class name is “Moon Phase”. In Figure 7(b), the class name is “Clock Time”.
Penguin Benchmark. The challenging Penguin dataset Penguin Research (2016) consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. (2016). The test set of this dataset is quite large, containing thousands of images with penguin counts ranging from 0 to 213. To build a manageable benchmark, we randomly sample 100 images while preserving a balanced ground truth range. We refer the reader to Appendix D for more details about this benchmark.
@@ 4.2 Results
The following section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section 4.1. For visual examples of LVLM-Count’s performance on these benchmarks, see Appendix B. The ablation study results are discussed in Appendix A. Additionally, Appendix K provides the performance of SOTA counting models on these benchmarks as a reference for evaluating the LVLMs’ performance. Also, an inference time analysis is provided in Appendix J.
@@
FSC-147. We evaluate the performance of various LVLMs and their enhanced performance using our method, LVLM-Count, on the test set of the FSC-147 dataset. We consider two baselines: (i) the baseline where the number of target segmentation masks is taken as the final answer, and (ii) the baseline where an image of the generated segmentation masks is provided to GPT-4o with a prompt to count the number of masks. Our experiments involve GPT-4o (a leading proprietary model), as well as two open-source models: Qwen2 VL 72B AWQ (Yang et al., 2024) and Gemma 3 27B Team et al. (2025). The expression $E$ used in different stages of our method corresponds to the category name provided in the test set. A simple query $Q$ in the form of “How many $E$ are there?” is constructed and used as the text prompt for the LVLM during the counting stage. In all experiments, the detection thresholds are set to $0.1$. The results are presented in Table 1, reporting the mean absolute error (MAE) and root mean square error (RMSE). Our findings indicate that LVLM-Count improves the performance of all three LVLMs in terms of MAE. Interestingly, while the base Qwen2 and Gemma 3 models are much less powerful than their commercial counterpart, they outperform the base GPT-4o when integrated into our pipeline.
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of target segmentaion masks 44.14 - 154.39 - GPT-4o counting target segmentaion masks 38.45 - 156.11 - GPT-4o 25.17 - 137.86 - LVLM-Count (GPT-4o as LVLM) 18.09 $\downarrow 7.08$ 88.77 $\downarrow 49.09$ Gemma 3 27B 30.23 - 136.43 - LVLM-Count (Gemma 3 27B as LVLM) 19.70 $\downarrow 10.53$ 110.62 $\downarrow 25.81$ Qwen2 VL 72B AWQ 33.45 - 145.90 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.67 $\downarrow 10.78$ 124.73 $\downarrow 21.17$
@@
Table 1: Evaluation on the test set of the FSC-147 dataset. In all tables, the results for base LVLMs and LVLM-Count are reported over three trials. Also, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation. To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix I. Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix H.
PASCAL VOC Benchmark. We evaluate the performance of three LVLMs—GPT-4o, Qwen2, and Gemma 3—on this benchmark, both with and without LVLM-Count. As shown in Table 2, LVLM-Count consistently outperforms the base LVLMs.
@@
Method MAE $\downarrow$ $\Delta$ RMSE $\downarrow$ $\Delta$ Number of the target segmentation masks 4.03 - 7.47 GPT-4o counting target segmentaion masks 6.60 - 15.09 GPT4o 4.46 - 8.35 - LVLM-Count (GPT4o as LVLM) 3.31 $\downarrow$ 1.15 7.11 $\downarrow$ 1.24 Gemma 3 27B 3.37 - 7.49 - LVLM-Count (Gemma 3 27B as LVLM) 2.87 $\downarrow$ 0.5 5.84 $\downarrow$ 1.65 Qwen2 VL 72B AWQ 4.83 - 8.84 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 4.05 $\downarrow$ 0.78 7.72 $\downarrow$ 1.12

### v4

@@ Abstract
Counting is a fundamental operation for various real-world visual tasks, requiring both object recognition and robust counting capabilities. Despite their advanced visual perception, large vision-language models (LVLMs) are known to struggle with counting tasks. In this work, we evaluate the performance of several LVLMs on visual counting tasks across multiple counting and vision datasets. We observe that while their performance may be less prone to error for small numbers of objects, they exhibit significant weaknesses as the number of objects increases. To alleviate this issue, we propose a simple yet effective baseline method that enhances LVLMs’ counting ability for large numbers of objects using a divide-and-conquer approach. Our method decomposes counting problems into sub-tasks. Moreover, it incorporates a mechanism to prevent objects from being split during division, which could otherwise lead to repetitive counting—a common issue in a naive divide-and-conquer implementation. We demonstrate the effectiveness of this approach across various datasets and benchmarks, establishing it as a valuable reference for evaluating future solutions.
@@ 1 Introduction
These applications often require distinguishing between objects of the same class with subtle variations, as well as recognizing complex concepts. Models trained solely on counting datasets struggle to generalize to such scenarios due to the limited availability of annotated data for fine-grained counting. Recent advancements in large vision-language models (LVLMs), such as GPT-4o (Achiam et al., 2023), combined with the massive scale of web-scraped training data, have enabled unprecedented zero-shot recognition capabilities, making them a promising candidate for handling complex and fine-grained counting tasks. However, evaluations of LVLMs have also revealed notable weaknesses in their numerical reasoning (Yin et al., 2023; Xu et al., 2024; Yang et al., 2023).
In this work, we focus specifically on the visual counting, one of the most fundamental aspects of numerical reasoning. We observe that although LVLMs perform well in counting small numbers of objects—typically fewer than 20—their accuracy deteriorates with larger quantities. Inspired by prior work on the rapid and accurate estimation of small quantities by Chattopadhyay et al. (2017) we propose a simple yet effective baseline method to alleviate this issue. Leveraging a divide-and-conquer approach, we engineer a simple pipeline that divides an image into carefully cut sub-images, and prompts the LVLM to count the objects of interest in each sub-image. The counts from the sub-images are then aggregated to make the final prediction. A key feature of our pipeline is a mechanism that prevents objects of interest from being split by the dividing lines, which could otherwise lead to double-counting. To have this mechanism two existing pre-trained detection and segmentation models are employed in the pipeline. The workflow is illustrated in Figure˜1.
@@
    We propose a solution for object-aware division. Accurate division is crucial, as parts of cut objects can lead to over-counting (see Figure˜2(a)). The proposed solution divides images without cutting through objects of interest specified by an arbitrary prompt.
@@
Figure 2: Comparison of the naive and the object-aware division. The objects of interest are the circles. In Figure˜2(a), we illustrate a naive division of the input image, which is divided into equally sized sub-images with straight lines. In Figure˜2(b), we illustrate the object-aware division, which avoids cutting through circles. In Figure˜2(c), we illustrate the counting error of GPT-4o for images with randomly positioned circles. The absolute counting error is the absolute difference between the ground truth and the number predicted by GPT-4o. The results are averaged over three trials.
@@ 2 Related Work
Text-based counting models. With the advent of vision-language foundation models such as CLIP and GroundingDINO, text-based open-world models have been proposed that are trained specifically for counting. Leveraging the rich textual and visual feature extraction capabilities of foundation models, obtained through web-scale training, the text-based counting models by Amini-Naieni et al. (2023); Dai et al. (2024); Kang et al. (2024); Amini-Naieni et al. (2024) have started to demonstrate comparable or superior accuracy. In addition, Shi et al. (2024) introduce TFOC, a counting model that does not require any counting-specific training. Instead, they cast the counting problem as a prompt-based segmentation task, using SAM (Kirillov et al., 2023) to obtain segmentation masks that determine the output number. Despite this progress, these models remain constrained by two principal limitations. First, their performance can degrade significantly under considerable distribution shift in the input samples. Second, the prompts they interpret are typically limited to simple object categories or referring expressions; as task complexity increases, these models often fail. Employing LVLMs for counting offers a potential pathway to mitigate these issues.
@@ 3.1 Area Detection
In this part of the pipeline, we assume that we are given a counting question $Q$ along with an image. The question $Q$ contains an expression $E$ that specifies a set of objects of interest. The expression $E$ distinguishes these objects from objects of other categories or the same category but with different attributes present in the image. By employing an LLM, the expression $E$ is extracted from $Q$. For example, let $Q$ be “How many brown eggs are in the image?”. $Q$ is given to an LLM, which is prompted to return the expression $E$, “brown eggs”, referring to the objects we want to count. After $E$ is extracted, it is provided as input to GroundingDINO along with the image. The output of GroundingDINO may be a single bounding box or a set of bounding boxes that have relevance to $E$ beyond a certain threshold. These bounding boxes often overlap and typically contain repeated objects. Thus, all the overlapping output bounding boxes are merged. After merging, a set of non-overlapping areas of interest may remain. We consider the non-overlapping areas as “detected areas”, which are then cropped to be passed to the next stage. This process is illustrated in Figure˜3.
@@ 3.2 Target Segmentation
The cropped images from the first stage contain objects of interest, and the ultimate goal is to divide them without cutting through those objects. However, a prerequisite for implementing such a mechanism is to first detect and localize the objects of interest. Each cropped image is fed into an open-world detection model along with $E$. The output of the open-world detection model produces a bounding box for each object of interest. The bounding boxes are then given as input to a segmentation model, which returns segmentation masks for the objects within each bounding box. We illustrate an example of this process in Figure˜4.
@@
Figure 4: Illustration of the target segmentation step of LVLM-Count. The goal is to produce all the instance masks for $E$ set to “brown egg”. The cropped detected area from Figure˜3, together with $E$, is given as input to GroundingDINO, which produces the output shown in Figure˜4(a). Figure˜4(a) is then given as input to SAM, which produces the output shown in Figure˜4(b).
@@
How to determine the segmentation masks. We employ a pre-trained segmentation model, specifically SAM (Kirillov et al., 2023), for the segmentation task. This model accepts bounding boxes as prompts and generates masks covering the most prominent objects within these boxes. However, we do not use SAM’s output masks directly, as in crowded scenes or cases with multiple occlusions, the masks often overlap, making it difficult to identify reliable division paths. To address this issue, we apply several post-processing steps. First, we perform non-maximum suppression on SAM-generated masks to eliminate those with significant overlaps corresponding to less certain bounding boxes. Additionally, we apply an erosion function to the segmentations, ensuring adjacent masks maintain a minimum separation of two pixels. For a detailed analysis and visual examples of how these post-processing operations enhance robustness in crowded scenes and occlusion cases, see Appendix˜C. The final processed masks for each cropped area are then passed to the next stage of our pipeline.
Robustness to the Accuracy of Area Detection and Target Segmentation Stages. Although we employ GroundingDINO—a state-of-the-art model for grounding and detection—it lacks the flexibility and robustness of LVLMs when generalizing to unseen data and complex concepts. Consequently, we minimizes dependence on the accuracy of the detection model used for area detection and target segmentation. Rather than prioritizing accuracy, we emphasize: i) not missing any region containing objects of interest during area detection, and ii) not missing segmentation masks for any objects of interest during target segmentation. These objectives are easily achieved by setting GroundingDINO’s detection threshold to a very low value in both stages.
@@
We demonstrate the effectiveness and robustness of our strategy in the early pipeline stages by showing that LVLM-Count significantly improves LVLMs’ performance on one of the most challenging counting datasets—the Penguin Dataset Penguin Research (2016), as detailed in Section˜4.2. This dataset is particularly difficult due to frequent occlusion, camouflage, and complex backgrounds Arteta et al. (2016). Furthermore, in our ablation studies, we test a variant of our pipeline that removes the detection model entirely. Instead, SAM is configured to generate segmentation masks for all objects in the scene. As shown in Table˜5, our method still enhances counting performance in this configuration, underscoring its robustness and minimal dependence on the detection model during the initial pipeline stages.
@@ 3.3 Object-aware Division
How to determine the starting and ending points of the paths. We utilize an unsupervised and non-parametric approach, to obtain the start and end points of the paths. A few pixels are sampled from each of the masks. To determine the location of the division paths on the $x$-axis, the samples taken from the masks are projected onto the $x$-axis. The projected points are automatically clustered using a non-parametric mean-shift algorithm [^1] In our implementation, we employ the MeanShift algorithm from the Scikit-learn library (Pedregosa et al., 2011), utilizing its default heuristic for automatic bandwidth estimation. (Comaniciu and Meer, 2002). Once the clusters are identified, the point between the point with the highest $x$ value in one cluster and the point with the lowest $x$ value in the subsequent cluster is considered the $x$-coordinate of a division path. In effect, knowing the $x$-coordinate of a vertical path means that the coordinates of its endpoints are known. In particular, assuming height $h$ for an image crop, we consider $P_{s}=(x,0)$ and $P_{e}=(x,h)$ as the start and end points, respectively. Note that using this technique, we obtain the appropriate coordinates for the division paths, as well as the number of paths. For example, if there is only one cluster along the $x$-axis, no division is required, and if there are two clusters, one vertical path will divide the image into two parts. We illustrate this approach in Figure˜5.
@@
How to draw the paths. Previous step obtains the endpoints of the division paths. Assume $P_{s}=(x,0)$ and $P_{e}=(x,h)$ are the start and end points of a vertical division path, respectively. In an ideal case where there are no masks in the path of a straight line connecting the two points, this path will be drawn by connecting all the pixels on the straight line. However, there are potential masks that can be considered obstacles blocking the path. In other words, beginning from $P_{s}$, the line needs to go around these obstacles to reach $P_{e}$. Consequently, we treat this as a 2-dimensional path-finding problem. To solve the problem, we build a 2D binary map, $I_{B}$, where the pixels covered by the masks are turned into black, indicating them as obstacles, and all the other pixels are turned into white, showing they are open for passage. This binary image $I_{B}$ is mapped into a graph $G$, where each white pixel is a node, and it is connected to all of its white neighboring pixels. We use the $A^{*}(G,P_{s},P_{e},g)$ search algorithm to find a path that connects $P_{s}$ to $P_{e}$, where the heuristic $g$ is set to be Manhattan distance. The output of $A^{*}$ is a set of connected pixels that go around the obstacles and connect $P_{s}$ to $P_{e}$, creating an object-aware division path, as shown in Figure˜6. The path-finding algorithm is run for all division coordinates. Finally, we draw the image contours based on these division paths and take the area surrounded by each contour as a resulting sub-image. Note that although not part of the main LVLM-Count pipeline, the above two steps can be applied to the $y$-axis in the same manner to obtain horizontal division paths. For further discussion, the reader may refer to Appendix˜S.
@@ 4 Experiments
In this section, we present the performance results of different LVLMs and their enhanced performance results using our method on a counting-specific dataset, a counting benchmark taken from a popular vision datasets, and a challenging counting benchmark that we propose using emoji icons. Additionally, we show the success of our design in enhancing counting performance on one of the most challenging counting datasets that features heavy occlusions and complex backgrounds. The code to reproduce the experiments is available at our GitHub repository  [^2] Our GitHub repository is available at: https://github.com/MuhammadFetrat/lvlm-count.
@@ 4.1 Datasets and Benchmarks
Emoji-Count. To our knowledge, no counting benchmark exists with large number of objects in a scene involving complex reasoning. To this end, we propose a challenging counting benchmark using emoji icons. From the $1816$ standard emoji icons, we remove those that directly overlap with concepts demonstrated by other icons. We then group the remaining $1197$ icons into $82$ classes. In each class, there are icons from the same or similar object categories, but with subtle differences that require complex reasoning to distinguish. For each of the $82$ classes, an empty $1024\times 1024$ image is first created. This image is filled with up to six categories chosen randomly from the class, with each category having a random count between $30$ and $50$ in the image. For each image, we create questions that ask the number of instances of the available categories in the image. This results in 415 image-question pairs. We illustrate two examples of this dataset in Figure˜7.
@@
Figure 7: Illustration of a smaller version of two challenging cases from Emoji-Count. In Figure˜7(a), the class name is “Moon Phase”. In Figure˜7(b), the class name is “Clock Time”.
Penguin Benchmark. The challenging Penguin dataset Penguin Research (2016) consistently exhibits heavy occlusion and complex background patterns that can easily be mistaken for penguins Arteta et al. (2016). The test set of this dataset is quite large, containing thousands of images with penguin counts ranging from 0 to 213. To build a manageable benchmark, we randomly sample 100 images while preserving a balanced ground truth range. We refer the reader to Appendix˜D for more details about this benchmark.
@@ 4.2 Results
This section presents the numerical results of our experiments with the base LVLMs and their corresponding LVLM-Count on each benchmark described in Section˜4.1. In the experiments, we consider two baselines: (i) a baseline where the number of target segmentation masks is taken as the final answer, and (ii) a baseline where an image of the generated segmentation masks is provided to GPT-4o along with a prompt to count them. For reference, we also report the performance of three state-of-the-art (SOTA) text-based counting-specific models: GroundingREC (Dai et al., 2024), CountGD (Amini-Naieni et al., 2024), and $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024).
It is important to note that the primary aim of this work is to propose a method that enhances the counting ability of an arbitrary large vision-language model (LVLM) across diverse datasets. The inclusion of these specialized models in our comparison should not be interpreted as an attempt to replace or compete with them on specific counting tasks—particularly those involving simple counting. These models are trained on dedicated counting datasets and perform efficiently on test samples from the same distribution or datasets with high category overlap. However, their performance may degrade significantly when applied to out-of-distribution samples. Furthermore, these models are constrained by the type and complexity of the input prompts they can process; they typically only accept simple category names or referring expressions, and often fail when presented with complex phrases or sentences.
In contrast, our method is designed to be comprehensive. Although it may not match the efficiency of task-specific models on datasets they were trained on, it performs robustly across a wide range of test distributions, despite not being explicitly trained on any of the tasks. Moreover, it is capable of handling arbitrary prompt types and structures, regardless of complexity. We also report the performance of TFOC (Shi et al., 2024), a SOTA training-free model. In essence, TFOC is a segmentation-based counting approach, similar to our baselines, but employs a more sophisticated methodology.
The above discussion is reflected in our experimental results. For the trained counting models, we report performance using weights obtained from training on the FSC-147 training set. These models perform well on the FSC-147 dataset and on the PASCAL VOC benchmark, which has high category overlap with FSC-147. However, on the Emoji-Count dataset—which exhibits a different distribution and requires complex understanding—their performance drops considerably. In contrast, our method, LVLM-Count, maintains high performance even on challenging benchmarks such as Emoji-Count. Another experiment on the TallyQA-Complex benchmark, which contains real-world scenes with complex counting questions, further reveals the limited complex reasoning capability of conventional trained models. LVLMs and our proposed LVLM-Count, however, demonstrate a superior ability to handle such challenging cases. The corresponding results on the TallyQA-Complex benchmark are provided in Appendix˜T.
We also conduct an ablation study on the effectiveness of different components of our pipeline in Appendix˜A. In summary, the ablation experiments support three conclusions: (i) each component has a positive effect over the baseline LVLM, and when combined, the full pipeline in Figure˜1 achieves the best performance compared to different variants; (ii) naive division with straight lines cannot replace object-aware division; and (iii) LVLM-Count is robust to the accuracy of the target segmentation stage. For visual examples of LVLM-Count’s performance on the benchmarks introduced in Section˜4.1, see Appendix˜B. An inference time analysis is provided in Appendix˜J. This analysis shows that the largest portion of the time is spent querying the LVLM, which is a step common to both the base LVLM and LVLM-Count. The other steps contribute a much smaller share to the overall inference time.
@@
FSC-147. We evaluate the performance of various LVLMs and their enhanced performance using our method, LVLM-Count, on the test set of the FSC-147 dataset. We consider two baselines: (i) the baseline where the number of target segmentation masks is taken as the final answer, and (ii) the baseline where an image of the generated segmentation masks is provided to GPT-4o with a prompt to count the number of masks. Our experiments involve GPT-4o (a leading proprietary model), as well as two open-source models: Qwen2 VL 72B AWQ (Yang et al., 2024) and Gemma 3 27B Team et al. (2025). The expression $E$ used in different stages of our method corresponds to the category name provided in the test set. A simple query $Q$ in the form of “How many $E$ are there?” is constructed and used as the text prompt for the LVLM during the counting stage. In all experiments, the detection thresholds are set to $0.1$. The results are presented in Table˜1, reporting the mean absolute error (MAE) and root mean square error (RMSE). Our findings indicate that LVLM-Count improves the performance of all three LVLMs in terms of MAE. Interestingly, while the base Qwen2 and Gemma 3 models are much less powerful than their commercial counterpart, they outperform the base GPT-4o when integrated into our pipeline.
@@
Table 1: Evaluation on the test set of the FSC-147 dataset. In all the tables of this section, the results for base LVLMs and LVLM-Count are reported over six trials. In the tables, columns marked with $\Delta$ show the performance difference between the LVLM-Count and the base LVLM it uses. Green indicates improvement, while red represents degradation. Also, the 95% confidence intervals for the MAEs are reported in the column named MAE 95% CI. To see the measured accuracy metrics for this dataset and the subsequent benchmarks, refer to Appendix˜I. Additionally, MAE analysis across different intervals of ground truth values for FSC-147 is provided in Appendix˜H.
@@
Method Trained Model MAE $\downarrow$ $\Delta$ MAE 95% CI RMSE $\downarrow$ $\Delta$ TFOC (Shi et al., 2024) 24.79 - - 137.15 - $\text{DAVE}_{\text{prm}}$ (Pelhan et al., 2024) 14.90 - - 103.42 - CountGD (Amini-Naieni et al., 2024) 14.76 - - 120.42 - GroundingREC (Dai et al., 2024) 10.12 - - 107.19 - Number of target segmentaion masks 44.14 - - 154.39 - GPT-4o counting target segmentaion masks 38.45 - - 156.11 - GPT-4o 25.57 - [24.74, 26.39] 137.26 - LVLM-Count (GPT-4o as LVLM) 17.86 $\downarrow 7.71$ [16.96, 18.77] 91.71 $\downarrow 45.55$ Gemma 3 27B 30.59 - [30.22, 30.97] 132.61 - LVLM-Count (Gemma 3 27B as LVLM) 20.25 $\downarrow 10.34$ [19.62, 20.89] 110.45 $\downarrow 22.16$ Qwen2 VL 72B AWQ 34.18 - [32.85, 35.51] 149.49 - LVLM-Count (Qwen2 VL 72B AWQ as LVLM) 22.29 $\downarrow 11.89$ [21.63, 22.96] 119.46 $\downarrow 30.44$
PASCAL VOC Benchmark. We evaluate the performance of three LVLMs—GPT-4o, Qwen2, and Gemma 3—on this benchmark, both with and without LVLM-Count. As shown in Table˜2, LVLM-Count consistently outperforms the base LVLMs.